# Signatures in TQFT : Asymptotics and Modularity

### v1

@@ 1. Introduction
The proof of this theorem involves several steps and occupies Sections 3 and 4. We first apply in Section 3 a finite Fourier transform in Equation (1) to transform this expression into a sum of evaluations of a rational function at roots of unity indexed by $\Delta_{p}$. We then apply Brion’s theorem to get an explicit trigonometric sum computing $\sigma_{2}(q/p)$ (see Theorem 3.1 for the precise statement): this step already presents miraculous cancellations for which we do not have any conceptual explanation.
Then in Section 4, we analyse the asymptotic behavior of the trigonometric sum in the most technical part of the paper. In order to control its denominators, we need to impose that the coefficients of the continued fraction expansion of $\theta$ do not grow too quickly. We use for that some classical results of Khinchin on metric Diophantine approximation and a variation on a technique due to Zaremba.
@@
This functional equation suggests that a similar integral version holds for signatures as stated in the following conjecture, checked by computer for all $0<q<p<100$.
@@
$\sigma_{2}\Big(\frac{q}{2q+p}\Big)-\sigma_{2}\Big(\frac{q}{p}\Big)=2q^{2}+2qp+p^{2}-1.$
This formula is reminiscent of the reciprocity formula of Dedekind sums: in fact we will make this analogy more precise in Section 6.1. We prove that the signature at $\zeta=e^{i\pi q/p}$ of the TQFT-vector space associated to the 4 times punctured sphere (with colors (p1)/2(p-1)/2 at the marked points) is actually equal to a 2-smoothed version, $S(q/p)$, of the Dedekind sum $s(q,p)$. It is then classical that this function $S(q/p)$ is the boundary value of a modular form (this time of weight $1/2$) and that it satisfies the following transformation law
@@
We provide an elementary proof of these facts for the benefit of the non-specialist reader.
This very satisfactory state of things would hopefully generalize to higher genus as well. However, even in genus 2, a precise (non asymptotic) relation between the signature $\sigma_{g}(q/p)$ and the Eichler integral is still lacking. Such a relation would presumably lead to a proof of the conjectured transformation law for $\sigma_{2}$.
@@
Acknowledgements: It is our pleasure to thank Pierre Charollois for early discussions on many aspects of this paper, Gaëtan Chenevier for his expertise on modular forms, and Don Zagier for very helpful explanations about Eichler integrals and quantum modular forms.
@@ 3. A trigonometric formula for the signature in genus 2
$\sigma_{2}({{\textstyle\frac{q}{p}}})=\frac{1-p^{2}}{6p^{2}}+\frac{1}{4p^{2}}\sum_{n=1,\,\mathrm{odd}}^{p-2}\frac{f(n,p,q)}{\sin^{3}(\frac{n\pi}{2p})\sin^{2}(\frac{qn\pi}{p})}$
@@
$=\frac{(x^{3}+3x^{-1})p\,\varphi(X)-(3x+x^{-3})p\,\varphi(X^{-1})}{2(x-x^{-1})^{3}(X-X^{-1})^{2}}$
@@ 4. Asymptotic behavior of the signature in genus 22
Our proof of this theorem makes heavy use of special properties of convergents, which will be reviewed below. We do not know the answer to the following question:
@@
We remark that for almost all irrational $\theta$, the convergents $\frac{q_{k}}{p_{k}}$ have the property that both $q_{k}$ and $p_{k}$ are odd for infinitely many $k$ (see Remark 4.7 below). But in Theorem 4.1, we don’t need to restrict to the convergents where both $q_{k}$ and $p_{k}$ are odd. We can take $\sigma_{2}(\frac{q}{p})$ to be the function defined by the expression in Theorem 3.1, so that $\sigma_{2}(\frac{q_{k}}{p_{k}})$ makes sense also when one of $q_{k}$ or $p_{k}$ is even. Our proof will show that with this interpretation of $\sigma_{2}(\frac{q}{p})$ the theorem still holds.
Before embarking on the proof of Theorem 4.1, let us discuss the convergence of the sum on the right hand side. We recall that for any real number $\theta$, its irrationality exponent $\mu(\theta)$ is the supremum of the set of $\mu\in\mathbb{R}$ such that the inequality
@@
is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $p>0$.
It is well-known that almost all irrational numbers are diophantine, that is, have irrationality exponent $\mu=2$. In Theorem 4.1, we can and will therefore assume $\theta$ is diophantine. Thus for all $\varepsilon>0$, the inequality $|\theta-\frac{q}{p}|\geq{p^{-2-\varepsilon}}$ holds for all but a finite number of pairs $(p,q)$. Observe that we do not need to impose the condition that $p$ and $q$ are coprime anymore.
Let us see how this implies convergence of the sum on the R.H.S. of Theorem 4.1. Given $n\in\mathbb{N}$, one derives a lower bound for $\sin(\pi n\theta)$ in the following way. First, observe the basic inequality
(13) $|\sin(x)|\geq\frac{2}{\pi}|x|\ \ \ \ \mathrm{for\ \ all}\ \ |x|\leq\frac{\pi}{2}$
which will be used many times in what follows. Now take $\ell\in\mathbb{Z}$ such that $|n\theta-\ell|\leq\frac{1}{2}$: for such an $\ell$ one has $|\sin(n\pi\theta)|=|\sin(\pi(n\theta-\ell))|\geq 2|n\theta-\ell|$ by (13) and hence
(14) $|\sin(n\pi\theta)|\geq 2{n^{-1-\varepsilon}}$
@@
Remark 4.3.
Actually, for the convergence of the sum over $n$ in Theorem 4.1, it would be sufficient to assume that $\mu(\theta)<3$. But in our proof of Theorem 4.1, we assume that $\mu(\theta)=2$ and, furthermore, that $\theta$ satisfies the conditions spelled out in Theorem 4.5 below. These conditions hold again for almost all irrational $\theta$.
@@
    For all $i\geq 1$, we have
    $\frac{1}{p_{i}(p_{i}+p_{i+1})}<\Big|\theta-\frac{q_{i}}{p_{i}}\Big|<\frac{1}{p_{i}p_{i+1}}<\frac{1}{p_{i}^{2}}~.$
  5. (v)
    [Lagrange] [^3] For a proof, see Perron [P, §15, Satz 2.17]. For all $1\leq i<k$, and any $n,\ell\in\mathbb{Z}$ with $1\leq n<p_{i+1}$, we have
    $\Big|n\frac{q_{k}}{p_{k}}-\ell\Big|\geq\Big|p_{i}\frac{q_{k}}{p_{k}}-q_{i}\Big|\geq\frac{1}{p_{i}+p_{i+1}}~.$
Not so well-known but crucial for our purposes is the following
Theorem 4.5.
For almost all irrational $\theta$, the coefficients $a_{k}$ and the denominators $p_{k}$ of the regular continued fraction expansion of $\theta$ satisfy the following:
  1. (i)
    (Khinchin [Kh1]) One has $a_{k}<k^{2}$ for all but finitely many $k$.
  2. (ii)
    (Khinchin [Kh2], Lévy [L]) The limit $\lim_{k\rightarrow\infty}p_{k}^{1/k}=L$ exists and is independent of $\theta$. Moreover $L>1$.
Remark 4.6.
  1. (i)
    Theorem 4.5(i) is a corollary of a more general result of Khinchin, see [Kh1, Theorem 30]), which says the following. Given any sequence of positive numbers $\phi(k)$, then for almost all irrational $\theta$:
    1. $\bullet$
      If $\sum\frac{1}{\phi(k)}<\infty$, then $a_{k}<\phi(k)$ for all but finitely many $k$.
    2. $\bullet$
      If $\sum\frac{1}{\phi(k)}=\infty$, then $a_{k}>\phi(k)$ for infinitely many $k$.
  2. (ii)
    The constant $L$ in Theorem 4.5(ii) is called Lévy’s constant. Lévy showed that $L=\exp(\pi^{2}/12\log 2)$. Since $L>1$, for almost all $\theta$, the denominators $p_{k}$ grow exponentially with $k$.
@@ 2. $\bullet$
We start from the trigonometric formula for $\sigma_{2}(\frac{q}{p})$ given in Theorem 3.1. The first observation is that the term $(1-p^{2})/(6p^{2})$ in this formula stays bounded as $p\rightarrow\infty$ and so does not contribute to the limit of $\sigma_{2}(\frac{q_{k}}{p_{k}})/p_{k}^{2}$. Thus it suffices to show
(17) $\lim_{k\rightarrow\infty}\frac{1}{4p_{k}^{4}}\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{f(n,p_{k},q_{k})}{\sin^{3}({n\pi}/{2p_{k}})\sin^{2}({q_{k}n\pi}/{p_{k}})}=\frac{16}{\pi^{3}}\sum_{n\geq 1,\textrm{odd}}\frac{1}{n^{3}\sin(n\pi\theta)}$
with $f(n,p,q)$ given as in Theorem 3.1.
The next observation is that since $\lim{q_{k}}/{p_{k}}=\theta$, we have
$\lim_{k\rightarrow\infty}\frac{f(n,p_{k},q_{k})}{p_{k}}=(3+1+1+3)\sin(n\pi\theta)=8\sin(n\pi\theta)~,$
and using $\lim_{x\rightarrow 0}(x/\sin x)=1$ in the denominator on the left, we see that the limit of the $n$-th summand on the left of (17) is precisely the $n$-th summand on the right. (Here we define the $n$-th summand on the left to be zero if $n>p_{k}-2$.) Unfortunately, none of the classical conditions for interchanging the limit and the sum apply here, so this is (by far) not enough for proving the theorem. Therefore a more careful analysis is needed.
Recall that $f(n,p,q)$ has four summands. Let us focus on the first summand; the proof is the same for the other three. Taking out the factor $(3p_{k}-3)/4p_{k}$ whose limit is $3/4$, we must prove
(18) $\lim_{k\rightarrow\infty}\frac{1}{p_{k}^{3}}\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{\sin((2q_{k}-1)n\pi/2p_{k})}{\sin^{3}({n\pi}/{2p_{k}})\sin^{2}({q_{k}n\pi}/{p_{k}})}=\frac{8}{\pi^{3}}\sum_{n\geq 1,\textrm{odd}}\frac{1}{n^{3}\sin(n\pi\theta)}~.$
Applying the addition formula $\sin(\alpha+\beta)=\sin\alpha\cos\beta+\cos\alpha\sin\beta$ to its numerator, the sum on the left (before taking limits) is equal to $S^{(1)}(q_{k},p_{k})-S^{(2)}(q_{k},p_{k})$, where
(19) $S^{(1)}(q_{k},p_{k})$ $=$ $\frac{1}{p_{k}^{3}}\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{\cos(n\pi/2p_{k})}{\sin^{3}({n\pi}/{2p_{k}})\sin({q_{k}n\pi}/{p_{k}})}$
(20) $S^{(2)}(q_{k},p_{k})$ $=$ $\frac{1}{p_{k}^{3}}\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{\cos(q_{k}n\pi/p_{k})}{\sin^{2}({n\pi}/{2p_{k}})\sin^{2}({q_{k}n\pi}/{p_{k}})}$
Therefore (18) follows from the following two lemmas.
Lemma 4.8.
$\lim\limits_{k\rightarrow\infty}S^{(1)}(q_{k},p_{k})=\frac{8}{\pi^{3}}\sum\limits_{n\geq 1,\mathrm{odd}}\frac{1}{n^{3}\sin(n\pi\theta)}~.$
Lemma 4.9.
$\lim\limits_{k\rightarrow\infty}S^{(2)}(q_{k},p_{k})=0~.$
Proof of Lemma 4.8.
We assume that $\mu(\theta)=2$. It implies (see (14)) that for every $\varepsilon>0$ there is $C_{\varepsilon}>0$ so that
(21) $|\sin(n\pi\theta)|\geq C_{\varepsilon}{n^{-1-\varepsilon}}$
holds for all $n\geq 1$.
Let $S^{(1)}(q_{k},p_{k},\theta)$ denote the sum $S^{(1)}(q_{k},p_{k})$ but with $\sin(n\pi\theta)$ in the denominator in place of $\sin({q_{k}n\pi}/{p_{k}})$. Using (13), we see that the $n$-th summand of $S^{(1)}(q_{k},p_{k},\theta)$ is bounded from above by $1/(n^{3}\sin(n\pi\theta))$. So we can apply the Lebesgue dominated convergence theorem to this sum. We get
$\lim_{k\rightarrow\infty}S^{(1)}(q_{k},p_{k},\theta)=\frac{8}{\pi^{3}}\sum_{n\geq 1,\textrm{odd}}\frac{1}{n^{3}\sin(n\pi\theta)}~.$
Thus to prove Lemma 4.8 it suffices to show that the difference $d_{k}=S^{(1)}(q_{k},p_{k})-S^{(1)}(q_{k},p_{k},\theta)$ goes to zero as $k\rightarrow\infty$. We have
$d_{k}=\frac{1}{p_{k}^{3}}\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{\cos(n\pi/2p_{k})}{\sin^{3}({n\pi}/{2p_{k}})}\Biggl(\frac{1}{\sin({q_{k}n\pi}/{p_{k}})}-\frac{1}{\sin(n\pi\theta)}\Biggr)~.$
@@
We can bound the denominators from below using (13), (21), and the trivial lower bound $|\sin({q_{k}n\pi}/{p_{k}})|\geq\sin(\pi/p_{k})\geq 2/p_{k}$. Thus for all $\varepsilon>0$ we have
(22) $|d_{k}|\leq\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{1}{n^{3}}\frac{p_{k}}{2}\frac{n^{1+\varepsilon}}{C_{\varepsilon}}\Big|\sin(n\pi\theta)-\sin(\frac{q_{k}n\pi}{p_{k}})\Big|~.$
@@
$<$ $\frac{n\pi}{p_{k}^{2}}=\frac{\sqrt{n}\sqrt{n}\pi}{p_{k}^{2}}$
(24) $\leq$ $\frac{\sqrt{n}\pi}{p_{k}^{3/2}}$

### v2

@@ 1. Introduction
The proof of this theorem involves several steps and occupies Sections 3 and 4. We first apply in Section 3 a finite Fourier transform in Equation (1) to transform this expression into a sum of evaluations of a rational function at roots of unity indexed by $\Delta_{p}$. We then apply Brion’s theorem to get an explicit trigonometric sum computing $\sigma_{2}(q/p)$; this step presents miraculous cancellations for which we do not have any conceptual explanation.
Then in Section 4, we provide the asymptotic analysis needed to prove the theorem. Here we use a simplification of our trigonometric formula which was pointed out to us by Y. Murakami [Mu]. This simplification makes the asymptotic analysis much easier than in our original proof.
@@
This functional equation suggests that a similar integral version holds for signatures. In the first version of this paper (of December 15, 2025), we made the following conjecture, which we had checked by computer for all $0<q<p<100$.
@@
(3) $\sigma_{2}\Big(\frac{q}{2q+p}\Big)-\sigma_{2}\Big(\frac{q}{p}\Big)=2q^{2}+2qp+p^{2}-1~.$
In early February 2026, we were informed by Y. Murakami [Mu] that he has obtained a proof of this conjecture starting from our trigonometric formula for $\sigma_{2}(\textstyle{\frac{q}{p}})$.
Our conjectured formula (3) is reminiscent of the reciprocity formula of Dedekind sums: in fact we will make this analogy more precise in Section 6.1. We prove that the signature at $\zeta=e^{i\pi q/p}$ of the TQFT-vector space associated to the 4 times punctured sphere (with colors (p1)/2(p-1)/2 at the marked points) is actually equal to a 2-smoothed version, $S(q/p)$, of the Dedekind sum $s(q,p)$. It is then classical that this function $S(q/p)$ is the boundary value of a modular form (this time of weight $1/2$) and that it satisfies the following transformation law
@@
We provide an elementary proof of these facts for the benefit of the non-specialist reader. This very satisfactory state of things where signatures have modular properties would hopefully generalize to higher genus as well.
@@
Acknowledgements: It is our pleasure to thank Pierre Charollois for early discussions on many aspects of this paper, Gaëtan Chenevier for his expertise on modular forms, and Don Zagier for very helpful explanations about Eichler integrals and quantum modular forms. We also thank Yuya Murakami for sharing a draft of his paper with us in early February 2026.
@@ 3. A trigonometric formula for the signature in genus 2
(5) $\sigma_{2}({{\textstyle\frac{q}{p}}})=\frac{1-p^{2}}{6p^{2}}+\frac{1}{4p^{2}}\sum_{n=1,\,\mathrm{odd}}^{p-2}\frac{f(n,p,q)}{\sin^{3}(\frac{n\pi}{2p})\sin^{2}(\frac{qn\pi}{p})}$
@@
We thank Y. Murakami for pointing out the following simplification of this formula:
Corollary 3.2 ([Mu, Proposition 7.1]).
For $0<q<p$ coprime odd integers, we have
(6) $\sigma_{2}({{\textstyle\frac{q}{p}}})=\frac{2}{p}\sum_{n=1,\,\mathrm{odd}}^{p-2}\frac{\cot^{3}(\pi n/2p)}{\sin(\pi nq/p)}$
@@
Applying standard addition formulas for the sine and cosine functions in $f(n,p,q)$ and regrouping terms, the expression (5) can be rewritten as
(7) $\sigma_{2}({{\textstyle\frac{q}{p}}})=\frac{1-p^{2}}{6p^{2}}+\frac{2}{p^{2}}\sum_{n=1,\,\mathrm{odd}}^{p-2}\frac{\cos(\pi qn/p)}{\sin^{2}(\pi qn/p)}+\frac{2}{p}\sum_{n=1,\,\mathrm{odd}}^{p-2}\frac{\cot^{3}(\pi n/2p)}{\sin(\pi nq/p)}~.$
But the first term and the first sum in (7) cancel thanks to the following trigonometric identity:
Lemma 3.3 ([Mu, Equation (7.1)]).
For any odd $p>0$, one has
$\sum_{m=1,\,\mathrm{odd}}^{p-2}\frac{\cos(\pi m/p)}{\sin^{2}(\pi m/p)}=\frac{p^{2}-1}{12}~.$
Proof of Theorem 3.1.
@@
(17) $=\frac{(x^{3}+3x^{-1})p\,\varphi(X)-(3x+x^{-3})p\,\varphi(X^{-1})}{2(x-x^{-1})^{3}(X-X^{-1})^{2}}$
@@
Remark 3.5.
Murakami’s rewriting (7) of our trigonometric formula can be obtained directly from the above by writing $\varphi(X)=-(X+X^{-1})+p(X-X^{-1})$ and rewriting the expression (17) as follows:
$\psi(n,n,n)$ $=\frac{-(x-x^{-1})^{3}p(X+X^{-1})+(x+x^{-1})^{3}p^{2}(X-X^{-1})}{2(x-x^{-1})^{3}(X-X^{-1})^{2}}$
$=-\frac{p}{2}\,\frac{X+X^{-1}}{(X-X^{-1})^{2}}+\frac{p^{2}}{2}\,\frac{(x+x^{-1})^{3}}{(x-x^{-1})^{3}}\,\frac{1}{X-X^{-1}}~.$
Remembering that $x=e^{\frac{i\pi n}{2p}}$ and $X=x^{2q}$ and plugging this into Eq. (15) proves Formula (7).
@@ 4. Asymptotic behavior of the signature in genus 22
We remark that for almost all irrational $\theta$, the convergents $\frac{q_{k}}{p_{k}}$ have the property that both $q_{k}$ and $p_{k}$ are odd for infinitely many $k$ (see Remark 4.4). But in Theorem 4.1, we don’t need to restrict to the convergents where both $q_{k}$ and $p_{k}$ are odd. We can take $\sigma_{2}(\frac{q}{p})$ to be the function defined by the expression in Corollary 3.2, so that $\sigma_{2}(\frac{q_{k}}{p_{k}})$ makes sense also when one of $q_{k}$ or $p_{k}$ is even. Our proof will show that with this interpretation of $\sigma_{2}(\frac{q}{p})$ the theorem holds without parity restrictions on the convergents.
We do not know the answer to the following question:
@@
Before proving Theorem 4.1, let us discuss the convergence of the infinite sum on the right hand side. We recall that for any irrational number $\theta$, its irrationality exponent $\mu(\theta)$ is the supremum of the set of $\mu\in\mathbb{R}$ such that the inequality
@@
is satisfied by an infinite number of coprime integer pairs $(p,q)$ with $p>0$. One always has $\mu(\theta)\geq 2$, as follows from the continued fraction expansion (see Proposition 4.3(iv) below.) Moreover it is well-known that almost all irrational numbers $\theta$ are diophantine, that is, have irrationality exponent $\mu(\theta)$ precisely equal to $2$. In order to prove Theorem 4.1, it will be sufficient to assume that $\mu(\theta)<3$. Assuming this from now on, take $0<\varepsilon<3-\mu(\theta)$, then the inequality $|\theta-\frac{q}{p}|\geq{p^{-3+\varepsilon}}$ holds for all but a finite number of pairs $(p,q)$. Observe that we do not need to impose the condition that $p$ and $q$ are coprime anymore.
Let us see how this implies convergence of the sum on the R.H.S. of Theorem 4.1. Given $n\in\mathbb{N}$, one derives a lower bound for $\sin(\pi n\theta)$ in the following way. First, observe the inequality
(18) $|\sin(\pi x)|\geq 2|x|\ \ \ \ \mathrm{for\ \ all}\ \ |x|\leq\textstyle{\frac{1}{2}}~.$
Now take $\ell\in\mathbb{Z}$ such that $|n\theta-\ell|\leq\frac{1}{2}$: for such an $\ell$ one has $|\sin(n\pi\theta)|=|\sin(\pi(n\theta-\ell))|\geq 2n|\theta-\frac{\ell}{n}|$ by (18) and hence
(19) $|\sin(n\pi\theta)|\geq 2{n^{-2+\varepsilon}}$
@@
    For all $k\geq 1$, we have
    $\Big|\theta-\frac{q_{k}}{p_{k}}\Big|<\frac{1}{p_{k}^{2}}~.$
@@
We start from Murakami’s simplification of our trigonometric formula for $\sigma_{2}(\frac{q}{p})$ in Corollary 3.2. Thus we must show
(22) $\lim_{k\rightarrow\infty}2\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{\cos^{3}(\pi n/2p_{k})}{p_{k}^{3}\sin^{3}(\pi n/2p_{k})\sin(\pi nq_{k}/p_{k})}=\frac{16}{\pi^{3}}\sum_{n\geq 1,\textrm{odd}}\frac{1}{n^{3}\sin(n\pi\theta)}~.$
Since $\lim{q_{k}}/{p_{k}}=\theta$ and $\lim_{x\rightarrow 0}(x/\sin x)=1$, the limit as $k\rightarrow\infty$ of the $n$-th summand on the left of (22) is precisely the $n$-th summand on the right. (Here we define the $n$-th summand on the left to be zero if $n>p_{k}-2$.) Thus it suffices to show that when $\theta$ is irrational with $\mu(\theta)<3$, we may interchange the limit and the sum.
Let $S_{k}$ denote the sum on the L.H.S. of (22) (before taking the limit as $k\rightarrow\infty$) and let $S^{(\theta)}_{k}$ the same sum but with $\sin(n\pi\theta)$ in the denominator in place of $\sin({q_{k}n\pi}/{p_{k}})$. Using (18), we see that the $n$-th summand of $S^{(\theta)}_{k}$ is bounded from above by $16/(n^{3}\sin(n\pi\theta))$ which, as we have seen, for $\mu(\theta)<3$ is the general term of an absolutely convergent sum. Thus for $S^{(\theta)}_{k}$ we may interchange the limit and the sum, and we get
$\lim_{k\rightarrow\infty}S^{(\theta)}_{k}=\frac{16}{\pi^{3}}\sum_{n\geq 1,\textrm{odd}}\frac{1}{n^{3}\sin(n\pi\theta)}~.$
It remains to see that the difference $d_{k}=S_{k}-S^{(\theta)}_{k}$ goes to zero as $k\rightarrow\infty$. We proceed as follows. Note first that (19) implies that for every $0<\varepsilon<3-\mu(\theta)$ there is $C_{\varepsilon}>0$ so that
(23) $|\sin(n\pi\theta)|\geq C_{\varepsilon}{n^{-2+\varepsilon}}$
holds for all $n\geq 1$. We have
$d_{k}=\ 2\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{\cos^{3}(n\pi/2p_{k})}{p_{k}^{3}\sin^{3}({n\pi}/{2p_{k}})}\Biggl(\frac{1}{\sin({q_{k}n\pi}/{p_{k}})}-\frac{1}{\sin(n\pi\theta)}\Biggr)~.$
@@
We can bound the denominators from below using (18), (23), and the trivial lower bound $|\sin({q_{k}n\pi}/{p_{k}})|\geq\sin(\pi/p_{k})\geq 2/p_{k}$. Thus for $0<\varepsilon<3-\mu(\theta)$ we have
(24) $|d_{k}|\leq\ 2\sum_{n=1,\textrm{ odd}}^{p_{k}-2}\frac{1}{n^{3}}\frac{p_{k}}{2}\frac{n^{2-\varepsilon}}{C_{\varepsilon}}\Big|\sin(n\pi\theta)-\sin(\frac{q_{k}n\pi}{p_{k}})\Big|~.$
@@
$<$ $\frac{n\pi}{p_{k}^{2}}=\frac{n^{\varepsilon}n^{1-\varepsilon}\pi}{p_{k}^{1+\varepsilon}p_{k}^{1-\varepsilon}}$
(26) $\leq$ $\frac{n^{\varepsilon}\pi}{p_{k}^{1+\varepsilon}}$